Evaluate



lim 2r Σ r=1 [2n^2/(n+r)^3]


n→∞

1
Expert's answer
2022-06-13T14:41:03-0400

"\lim\limits_{n\rightarrow\infty}{\sum_{r=1}^{2n}\frac{2n^2}{\left(n+r\right)^3}}=\lim\limits_{n\rightarrow\infty}{\sum_{r=1}^{2n}\frac{2n^2}{{n^3\left(1+\frac{r}{n}\right)}^3}}=\lim\limits_{n\rightarrow\infty}{\frac{1}{n}\sum_{r=1}^{2n}\frac{2}{\left(1+\frac{r}{n}\right)^3}}="

"=2\int_{\lim\limits_{n\rightarrow\infty}{\frac{1}{n}}}^{2}\frac{dx}{\left(1+x\right)^3}=2\int_{0}^{2}\frac{d\left(1+x\right)}{\left(1+x\right)^3}="

"=2\left.\frac{\left(1+x\right)^{-2}}{-2}\right|\begin{matrix}2\\0\\\end{matrix}=\left.-\frac{1}{\left(1+x\right)^2}\right|\begin{matrix}2\\0\\\end{matrix}=-\frac{1}{9}+1=\frac{8}{9}"